OpenStudy (anonymous):

What is the slope of the line that passes through (-4, 5) and (2, -3)?

OpenStudy (anonymous):

The slope of a line can be represented by \[m=(y_2-y_1)/(x_2-x_1)\] For the first point, x1 = -4 and y1 = 5. For the second point x2 = 2 and y2 = -3. We substitute to get\[m=[(-3)-(5)]/[(2)-(-4)]=-8/6=-4/3\]